OpenStudy (kmullis6):

Is the square root of 6 rational or irrational? Please explain...

OpenStudy (anonymous):

root(6) = root(2 * 3) = root(2) * root(3) Roots of both 2 and 3 are irrational so their product is also irrational. so, root(6) is irrational.

OpenStudy (kmullis6):

Thanks! (:

OpenStudy (anonymous):

U r welcome... n thanks